OpenStudy (anonymous):

A bag contains 10 marbles—three red, three blue, and four green. What is the probability of pulling a red marble, not replacing it, and then pulling a blue marble? 9/100 1/10 6/19 3/5

OpenStudy (anonymous):

To start, what do you think is the probability of pulling a red marble?

OpenStudy (anonymous):

pulling a red marble is 3/10, then pulling a blue no replacement is 3/9, the probability is (3/10)*(3/9)=9/90=1/10

OpenStudy (anonymous):

thanks that helpped alot :)

OpenStudy (anonymous):

@kittyxie1 I think it's better to help him find the answer, than giving it to him immediately. The site's goal is to help people learn.

OpenStudy (anonymous):

she did help me she explained it to me and i had another question i did the same thinh on and got it right